OpenStudy (anonymous):

f f(x) = log3 (x + 1), what is f−^1 (2)? 1 8 10 27

OpenStudy (helder_edwin):

\[\large f(x)=\log_3(x+1) \] right?

OpenStudy (anonymous):

yes

OpenStudy (helder_edwin):

\[\large y=\log_3(x+1) \] \[\large 3^y=x+1 \] \[\large x=3^y-1 \] \[\large f^{-1}(x)=3^x-1 \]

OpenStudy (helder_edwin):

therefore \[\large f^{-1}(2)=3^2-1=9-1=8 \]

OpenStudy (anonymous):

ty :)

OpenStudy (helder_edwin):

u r welcome